An object thrown vertically up from the ground passes the height 5m twice in an interval of 10 sec. What is its time of flight

Grade:11

1 Answers

Abhinav
8 years ago
particle passes height 5mtwice in an interval of 10 sec .So it took 5 sec for reaching heighest point of flight.
so
velocity at 5 m height is
0= v- gt
0=v-10*5
v=50
so speed at start = 502= u2- 2*10*5
u= 51 m/s
time taken to reach 5m =.1 sec
so total time of flight = 10.2